Please help me solve this Word Problem:
Kelly had 12 dollars and spent x dollars. Ami had 6 dollars an collected x dollars. The two girls then had the same amount of money. Form an equation and solve it to find the value of x.

Answer by ptaylor(2198)

Our equation to solve is:
12-x=6+x add x to and subtract 6 from each side
12-x+x-6=x+x collect like terms
6=2x divide each side by 2
x=$3--------------------------------value of x
CK
12-3=6+3
9=9
Hope this helps---ptaylor